# Commutation of Smyth and Hoare Power Constructions in Well-filtered Dcpos

### v3

@@ 5 The double power well-filtered dcpo
$\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})={\uparrow}\{\bigcup\mathcal{C}:\mathcal{C}\in\mathcal{HH}(L),\forall\mathcal{K}\in cl(\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\}),\mathcal{C}\cap\mathcal{K}\neq\emptyset\}$.
For each $A\in\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})$, we have ${\downarrow}\{A\}\in\mathcal{HH}(L)$ and $A=\bigcup{\downarrow}\{A\}$. Suppose that there exists a $\mathcal{K}\in cl(\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\})$ such that ${\downarrow}\{A\}\cap\mathcal{K}=\emptyset$, then $\mathcal{K}\in\Box(\mathcal{H}(L)\setminus{\downarrow}\{A\})$. By the well-filteredness of $\mathcal{H}(L)$, we know $\Box(\mathcal{H}(L)\setminus{\downarrow}\{A\})$ is a Scoot open subset of $\mathcal{QH}(L)$. Therefore, $\Box(\mathcal{H}(L)\setminus{\downarrow}\{A\})\cap\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\}\neq\emptyset$. So there is an $\mathcal{A}\in\mathbf{A}$ such that $\phi_{L}(\mathcal{A})\subseteq\mathcal{H}(L)\setminus{\downarrow}\{A\}$. This further indicates that $\phi_{L}(\mathcal{A})\cap{\downarrow}\{A\}=\emptyset$, i.e., $A\notin\phi_{L}(\mathcal{A})$. From which we derive that $A\cap K=\emptyset$ for some $K\in\mathcal{A}$. Note that $\mathcal{A}\in\mathbf{A}$ implies $K\in\bigcup\mathbf{A}$. As a result, clearly, $A\cap K=\emptyset$ contradicts the fact that $A\in\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})$. So $A=\bigcup{\downarrow}\{A\}\in\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})$, and hence, $\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})\subseteq\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})$.
Pick an arbitrary element $\bigcup\mathcal{C}$, where $\mathcal{C}\in\mathcal{HH}(L)$ and for each $\mathcal{K}\in cl(\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\}),\mathcal{C}\cap\mathcal{K}\neq\emptyset$. Assume that there exists a $K\in\bigcup\mathbf{A}$, i.e., $K\in\mathcal{A}$ for some $\mathcal{A}\in\mathbf{A}$, such that $\bigcup\mathcal{C}\cap K=\emptyset$. Then for each $C\in\mathcal{C}$, $C\cap K=\emptyset$, which immediately means that $C\notin\phi_{L}(\mathcal{A}),\forall C\in\mathcal{C}$, in other words, $\mathcal{C}\cap\phi_{L}(\mathcal{A})=\emptyset$. This contradicts $\bigcup\mathcal{C}\in\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})$. So $\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})\subseteq\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})$.
@@
From the above results, we derive that $\mathcal{Q}\mathcal{H}$ is a monad on $\mathbf{WF}$. Notably, Steven Vickers studied the Eilenberg-Moore algebras of the double powerlocale $P_{U}P_{L}$ (where $P_{U}$ and $P_{L}$ denote the upper and lower powerlocales, which commute up to isomorphisms via a distributive law) and it is concluded that $P_{U}P_{L}$-algebras correspond to localic frames [24]. Motivated by this work, we now investigate the algebraic structures of $\mathcal{Q}\mathcal{H}$-algebras.
We first give the unit $\gamma$ and the multiplication $\rho$ of $\mathcal{Q}\mathcal{H}$ with the following diagrams, which are closely related to that of $\mathcal{Q}$ and $\mathcal{H}$ according to Beck [2].
@@
([8]) Let $\mathbf{C}$ be a poset-enriched category. A monad $\mathcal{T}=(\mathcal{T},\eta,\mu)$ on $\mathbf{C}$ is called a right $\rm{KZ}$-monad if $\mathcal{T}$ is a poset-functor and satisfies $\eta_{TX}\leq T\eta_{X}$ for all $X\in\mathbf{C}$. Left $\rm{KZ}$-monads are defined poset-dually, by reversing the inequality.
@@
[8] Let $\mathcal{T}=(\mathcal{T},\eta,\mu)$ be a $\rm{KZ}$-monad on the category $\mathbf{C}$. Then by $(\mathrm{KZ}_{1})$, every object has at most one structure map.
@@
The category $\mathbf{WF}$ is a poset-enriched category; $\mathcal{H}$ and $\mathcal{Q}$ are both poset-functors.
@@
By Proposition 2.4 (1) and (2), $(\mathcal{Q}\mathcal{H}(L),\rho_{L})$ is a $\mathcal{Q}\mathcal{H}$-algebra and $\alpha:(\mathcal{Q}\mathcal{H}(L),\rho_{L})\rightarrow(L,\alpha)$ is a $\mathcal{Q}\mathcal{H}$-algebra homomorphism.
@@
    Theorem 5.3 and Proposition 5.5 imply that $\mathcal{Q}\eta:\mathcal{Q}\rightarrow\mathcal{Q}\mathcal{H}$ is a morphism of monads. By Lemma 5.4, $(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\mathcal{Q}\eta_{L})$ and $(L,\alpha\circ\mathcal{Q}\eta_{L})=(L,\alpha^{\mathcal{Q}})$ are $\mathcal{Q}$-algebras, and $\alpha:(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\mathcal{Q}\eta_{L})\rightarrow(L,\alpha^{\mathcal{Q}})$ is a $\mathcal{Q}$-algebra homomorphism. By Proposition 2.4 (1), $(\mathcal{Q}\mathcal{H}(L),\iota_{\mathcal{H}(L)})$ is also a $\mathcal{Q}$-algebra. Since $\mathcal{Q}$ is a right $\rm{KZ}$-monad, Proposition 5.8 indicates that structure maps for $\rm{KZ}$-monads are unique, $\rho_{L}\circ\mathcal{Q}\eta_{L}=\iota_{\mathcal{H}(L)}$ holds; therefore, $\alpha:(\mathcal{Q}\mathcal{H}(L),\iota_{\mathcal{H}(L)})\rightarrow(L,\alpha^{\mathcal{Q}})$ is a $\mathcal{Q}$-algebra homomorphism.
@@
The fact given in Proposition 5.11 that $L$ is a $\mathcal{Q}$-algebra and an $\mathcal{H}$-algebra yields that $L$ is indeed a meet continuous complete lattice by Corollary 3.8 and Proposition 3.16. Thus the proof will be completed if we prove that $L$ is distributive.
@@
Obviously, the union of arbitrary finite elements in $\mathcal{QH}(L)$ is still compact saturated. By Lemma 4.2, we know the complete lattice $\mathcal{H}(L)$ endowed with the Scott topology is coherent, that is, the arbitrary finite intersection of compact saturated subsets of $\mathcal{H}(L)$ is still in $\mathcal{QH}(L)$. Thus $\mathcal{QH}(L)$ is a lattice, in which the supremum and infimum of each finite subset are their union and intersection, respectively. This contributes to the distributivity of $\mathcal{QH}(L)$.
@@
Let $f:(L,\alpha)\rightarrow(M,\beta)$ be a $\mathcal{QH}$-algebra homomorphism. It was shown in Theorem 5.3 that $\mathcal{Q}\eta:\mathcal{Q}\rightarrow\mathcal{QH}$ and $\theta_{\mathcal{H}}:\mathcal{H}\rightarrow\mathcal{QH}$ are morphisms of monads; hence, by Lemma 5.4, we know $f:(L,\alpha\circ\mathcal{Q}\eta_{L})\rightarrow(M,\beta\circ\mathcal{Q}\eta_{M})$ is a $\mathcal{Q}$-algebra homomorphism and $f:(L,\alpha\circ\theta_{\mathcal{H}(L)})\rightarrow(M,\beta\circ\theta_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{H}(M)}})$ is an $\mathcal{H}$-algebra homomorphism. It follows that $f$ preserves finite infs and arbitrary sups, that is to say, $f$ is a frame homomorphism.
@@
See the deflationary semilattice in $\mathbf{WF}$ quoted in Example 3.17 $\mathcal{L}$ again, it has been proven not to be a $\mathcal{Q}$-algebra. Thus evidenced by Proposition 5.11, $\mathcal{L}$ is not a $\mathcal{QH}$-algebra, which indicates that $\mathcal{QH}$ cannot give the free deflationary semilattices for $\mathbf{WF}$.
@@
The general observation that the algebras of the double power construction are frames is due to S. Vickers [24] or perhaps P. Taylor [23], although Vickers works with locales and Taylor works in an abstract category of locally compact sober spaces. In an abstract category of locally compact sober spaces, Taylor proved that when the power constructions commute, the composite monad is equivalent to the monad defined as double exponentiation by the Sierpinski space, and characterized its algebras as frames. Furthermore, the characterizations work in more general categories when the necessary exponentials exist. We know every core-compact well-filtered dcpo is locally compact sober [17] and by Corollary 4.9, for each locally compact sober dcpo, the Hoare and the Smyth power constructions commute with each other. Thus, our composite monad $\mathcal{Q}\mathcal{H}$ is also equivalent to the monad defined as double exponentiation by the Sierpinski space in the core-compact case. Then a natural open question is whether this equivalence extends to a more general category of well-filtered dcpos satisfying conditions (i) and (ii) of Theorem 4.9. We leave this for future investigation.

### v5

@@ 5 The double power well-filtered dcpo
$\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})={\uparrow}\{\bigcup\mathcal{C}:\mathcal{C}\in\mathcal{HH}(L),\forall\mathcal{K}\in cl(\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\}),\mathcal{C}\bigcap\mathcal{K}\neq\emptyset\}$.
For each $A\in\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})$, we have ${\downarrow}\{A\}\in\mathcal{HH}(L)$ and $A=\bigcup{\downarrow}\{A\}$. Assume, for contradiction, that there exists a $\mathcal{K}\in cl(\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\})$ such that ${\downarrow}\{A\}\bigcap\mathcal{K}=\emptyset$. Then $\mathcal{K}\in\Box(\mathcal{H}(L)\setminus{\downarrow}\{A\})$. Since $\Sigma\mathcal{H}(L)$ is well-filtered, $\Box(\mathcal{H}(L)\setminus{\downarrow}\{A\})$ is a Scott open subset of $\mathcal{QH}(L)$. Therefore, $\Box(\mathcal{H}(L)\setminus{\downarrow}\{A\})\bigcap\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\}\neq\emptyset$. Hence, there exists an $\mathcal{A}\in\mathbf{A}$ such that $\phi_{L}(\mathcal{A})\subseteq\mathcal{H}(L)\setminus{\downarrow}\{A\}$, that is, $\phi_{L}(\mathcal{A})\bigcap{\downarrow}\{A\}=\emptyset$. Meanwhile, this indicates $A\notin\phi_{L}(\mathcal{A})$. It follows that $A\cap K=\emptyset$ for some $K\in\mathcal{A}$. Since $\mathcal{A}\in\mathbf{A}$ implies $K\in\bigcup\mathbf{A}$, the relation $A\cap K=\emptyset$ directly contradicts the fact that $A\in\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})$. So $A=\bigcup{\downarrow}\{A\}\in\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})$; consequently, $\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})\subseteq\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})$.
Pick an arbitrary element $\bigcup\mathcal{C}\in\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})$, where $\mathcal{C}\in\mathcal{HH}(L)$ and for every $\mathcal{K}\in cl(\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\}),\mathcal{C}\bigcap\mathcal{K}\neq\emptyset$. Assume that there exists a $K\in\bigcup\mathbf{A}$, i.e., $K\in\mathcal{A}$ for some $\mathcal{A}\in\mathbf{A}$, such that $\bigcup\mathcal{C}\cap K=\emptyset$. Then for each $C\in\mathcal{C}$, $C\cap K=\emptyset$. This immediately implies that $C\notin\phi_{L}(\mathcal{A})$ for all $C\in\mathcal{C}$; in other words, $\mathcal{C}\bigcap\phi_{L}(\mathcal{A})=\emptyset$. This contradicts the fact that $\bigcup\mathcal{C}\in\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})$. So $\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})\subseteq\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})$.
@@
From the above results, we derive that $\mathcal{Q}\mathcal{H}$ is a monad on $\mathbf{WF}$. Notably, S. Vickers studied the Eilenberg-Moore algebras of the double powerlocale $P_{U}P_{L}$ (where $P_{U}$ and $P_{L}$ denote the upper and lower powerlocales, which commute up to isomorphisms via a distributive law) and concluded that $P_{U}P_{L}$-algebras correspond to localic frames [24]. Motivated by this work, we now investigate the $\mathcal{Q}\mathcal{H}$-algebra structures.
We first give the unit $\gamma$ and the multiplication $\rho$ of $\mathcal{Q}\mathcal{H}$ with the following diagrams, which are closely related to that of $\mathcal{Q}$ and $\mathcal{H}$ according to J. Beck [2].
@@
([8]) Let $\mathbf{C}$ be a poset-enriched category. A monad $\mathcal{T}=(\mathcal{T},\eta,\mu)$ on $\mathbf{C}$ is called a right $\rm{KZ}$-monad if $\mathcal{T}$ is a poset-functor and satisfies $\eta_{\mathcal{T}X}\leq\mathcal{T}\eta_{X}$ for all $X\in\mathbf{C}$. Left $\rm{KZ}$-monads are defined poset-dually, by reversing the inequality.
@@
[8] Let $\mathcal{T}=(\mathcal{T},\eta,\mu)$ be a $\rm{KZ}$-monad on a category $\mathbf{C}$. Then by $(\mathrm{KZ}_{1})$, every object has at most one structure map.
@@
The category $\mathbf{WF}$ is a poset-enriched category; the functors $\mathcal{H}$ and $\mathcal{Q}$ are both poset-functors.
@@
By (1) and (2) of Proposition 2.4, we know $(\mathcal{Q}\mathcal{H}(L),\rho_{L})$ is a $\mathcal{Q}\mathcal{H}$-algebra and $\alpha:(\mathcal{Q}\mathcal{H}(L),\rho_{L})\rightarrow(L,\alpha)$ is a $\mathcal{Q}\mathcal{H}$-algebra homomorphism.
@@
    Theorem 5.3 and Proposition 5.5 imply that $\mathcal{Q}\eta:\mathcal{Q}\rightarrow\mathcal{Q}\mathcal{H}$ is a morphism of monads. By Lemma 5.4, $(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\mathcal{Q}\eta_{L})$ and $(L,\alpha\circ\mathcal{Q}\eta_{L})=(L,\alpha^{\mathcal{Q}})$ are $\mathcal{Q}$-algebras, and $\alpha:(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\mathcal{Q}\eta_{L})\rightarrow(L,\alpha^{\mathcal{Q}})$ is a $\mathcal{Q}$-algebra homomorphism. By Proposition 2.4 (1), $(\mathcal{Q}\mathcal{H}(L),\iota_{\mathcal{H}(L)})$ is also a $\mathcal{Q}$-algebra. Since $\mathcal{Q}$ is a right $\rm{KZ}$-monad, Proposition 5.8 indicates that structure maps for a $\rm{KZ}$-monad are unique, $\rho_{L}\circ\mathcal{Q}\eta_{L}=\iota_{\mathcal{H}(L)}$ holds; therefore, $\alpha:(\mathcal{Q}\mathcal{H}(L),\iota_{\mathcal{H}(L)})\rightarrow(L,\alpha^{\mathcal{Q}})$ is a $\mathcal{Q}$-algebra homomorphism.
@@
Since $L$ is both a $\mathcal{Q}$-algebra and an $\mathcal{H}$-algebra by Proposition 5.11, it follows from Corollary 3.8 and Proposition 3.16 that $L$ is a meet continuous complete lattice. Thus to complete the proof, it suffices to show that $L$ is distributive.
@@
Obviously, the union of arbitrary finite elements in $\mathcal{QH}(L)$ remains compact saturated. By Lemma 4.2, we know the complete lattice $\mathcal{H}(L)$ endowed with the Scott topology is coherent, that is, the intersection of any finite compact saturated subsets of $\mathcal{H}(L)$ is again a member of $\mathcal{QH}(L)$. Consequently, $\mathcal{QH}(L)$ forms a lattice, where the supremum and infimum of each finite subset are their intersection and union, respectively. This implies the distributivity of $\mathcal{QH}(L)$.
@@
Let $f:(L,\alpha)\rightarrow(M,\beta)$ be a $\mathcal{QH}$-algebra homomorphism. It is shown in Theorem 5.3 that $\mathcal{Q}\eta:\mathcal{Q}\rightarrow\mathcal{QH}$ and $\theta_{\mathcal{H}}:\mathcal{H}\rightarrow\mathcal{QH}$ are morphisms of monads; hence, by Lemma 5.4, $f$ induces a $\mathcal{Q}$-algebra homomorphism $f:(L,\alpha\circ\mathcal{Q}\eta_{L})\rightarrow(M,\beta\circ\mathcal{Q}\eta_{M})$ and an $\mathcal{H}$-algebra homomorphism $f:(L,\alpha\circ\theta_{\mathcal{H}(L)})\rightarrow(M,\beta\circ\theta_{\mathcal{H}(M)})$. Consequently, $f$ preserves finite infs and arbitrary sups; in other words, $f$ is a frame homomorphism.
@@
See again the deflationary semilattice $\mathcal{L}$ in $\mathbf{WF}$ quoted in Example 3.17, it has been proven not to be a $\mathcal{Q}$-algebra. Thus as evidenced by Proposition 5.11, $\mathcal{L}$ is not a $\mathcal{QH}$-algebra, which indicates that $\mathcal{QH}$ cannot give the free deflationary semilattices for $\mathbf{WF}$.
@@
The observation has been made by S. Vickers and also P. Taylor in a different context. S. Vickers has proved that the composite of the power constructions is equivalent to double exponentiation for exponentiable (i.e., locally compact) locales [24], and P. Taylor characterized the algebras of the double exponentiation monad on locally compact spaces as frames [23]. Furthermore, the characterizations work in more general categories when the necessary exponentials exist. We know every core-compact well-filtered dcpo is locally compact sober [17] and by Corollary 4.10, for each locally compact sober dcpo, the Hoare and the Smyth power constructions commute with each other. Thus, our composite monad $\mathcal{Q}\mathcal{H}$ is also equivalent to the monad defined as double exponentiation by the Sierpiński space in the core-compact case. Nevertheless, it remains a natural open question whether this equivalence can be extended to a broader category of well-filtered dcpos satisfying conditions (i) and (ii) of Theorem 4.9. We leave this for future work.